# LanPaint: Training-Free Diffusion Inpainting with Asymptotically Exact and Fast Conditional Sampling

### v2

@@ 3.1 Bidirectional Guided (BiG) Score
In RePaint and Langevin-based approaches, the inpainted region $\mathbf{x}_{t^{\prime}}$ is iterated towards high likelihood region of $p_{T-t^{\prime}}(\mathbf{x}\mid\mathbf{y})$, but $\mathbf{y}_{t^{\prime}}$ remains unaware of $\mathbf{x}_{t^{\prime}}$. This one-way dependency creates a critical flaw: if $\mathbf{x}_{t^{\prime}}$ enters a suboptimal region, $\mathbf{y}_{t^{\prime}}$ cannot receive corrective feedback, resulting in local maxima trapping of $\mathbf{x}_{t^{\prime}}$(Fig. 3). To escape such local optima, we propose jointly optimizing $\mathbf{x}_{t^{\prime}}$ and $\mathbf{y}_{t^{\prime}}$ through bidirectional feedback: while $\mathbf{x}_{t^{\prime}}$ is refined by $\mathbf{y}_{t^{\prime}}$ (as in prior work), $\mathbf{y}_{t^{\prime}}$ is also updated under the guidance of $\mathbf{x}_{t^{\prime}}$ while preserving observed content.
To achieve this, we observe that equation 7 can be relaxed to
$p_{t}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{o})\approx\frac{1}{Z}\,p_{t}(\mathbf{x}\mid\mathbf{y})\,\tfrac{p_{t}(\mathbf{y}\mid\mathbf{y}_{o})^{1+\lambda}}{p_{t}(\mathbf{y})^{\lambda}},$ (11)
with $\lambda>-1$ as the guidance scale and $Z$ is a normalizing constant. At $t=0$, the approximation ($\approx$) still becomes exact ($=$) at $t=0$. It is because $p_{0}(\mathbf{y}\mid\mathbf{y}_{o})=\delta(\mathbf{y}-\mathbf{y}_{o})$ is a delta distribution that remains invariant (up to normalization) when multiplied by other functions. Therefore $\tfrac{p_{t=0}(\mathbf{y}\mid\mathbf{y}_{o})^{1+\lambda}}{p_{t=0}(\mathbf{y})^{\lambda}}$ still represents $\delta(\mathbf{y}-\mathbf{y}_{o})$. The $\mathbf{x}$ component of its score function is still $\mathbf{s}_{\mathbf{x}}$ in equation 5, while the $\mathbf{y}$ component score can be approximated by the BIG score
@@
which is obtained by substituting $p_{t}(\mathbf{y})=\frac{p_{t}(\mathbf{x},\mathbf{y})}{p_{t}(\mathbf{x}|\mathbf{y})}$ into equation 11 then discard the unknown term $\nabla_{\mathbf{y}}\log p_{t}(\mathbf{x}|\mathbf{y})$. It successfully incorporated information of $\mathbf{x}_{t}$ as a guidance of the $\mathbf{y}_{t}$ sampling process.
The BiG score’s behavior depends on $\lambda$: when $\lambda=-1$, it reduces to unconditional sampling; at $\lambda=0$, it matches RePaint-like inpainting; and for $\lambda>0$, it enhances inpainting by penalizing unconditional scores $\mathbf{s}_{\mathbf{y}}$. Larger $\lambda$ values strengthen the corrective feedback from $\mathbf{x}_{t}$, helping $\mathbf{y}_{t}$ escape local optima more effectively.
@@
Though the unknown $\nabla_{\mathbf{y}}\log p_{t}(\mathbf{x}|\mathbf{y})$ term is discarded when deriving the BIG score, it still yields asymptotically exact conditional samples given partial observation, as this term is negligible near $t=0$ compared to the score of $p_{t}(\mathbf{y}|\mathbf{y}_{0})$, as the the following theorem states:
@@
Under the dynamics equation 13, the joint state $(\mathbf{x}_{t},\mathbf{y}_{t})$ converges to the distribution
$\mathbf{x}_{t},\,\mathbf{y}_{t}\;\sim\;\frac{1}{Z}\,p_{t}(\mathbf{x}\mid\mathbf{y})\,\tfrac{p_{t}(\mathbf{y}\mid\mathbf{y}_{o})^{1+\lambda}}{p(\mathbf{y}_{t})^{\lambda}}+o(\sqrt{1-\bar{\alpha}_{t}}),$ (14)
where $Z$ is a normalizing constant. Consequently, at $t=0$, the marginal $\mathbf{x}_{0}\sim p_{0}(\mathbf{x}\mid\mathbf{y}_{o})$ is an exact conditional sample, provided the Langevin dynamics converge. (Proof in Appendix E)
@@ 3.2 Fast Langevin Dynamics (FLD)
Solving Langevin dynamics (Eq. equation 13) is challenging: direct discretization requires trading step size against performance. Large steps accelerate convergence but introduce significant errors that yield noisy results, while small steps cause impractically slow convergence. We therefore seek an accelerated scheme with a stable solver, ensuring fast convergence to the stationary distribution while tolerating larger steps.
For fast convergence to the stationary distribution, existing approaches include Underdamped Langevin Dynamics (ULD) Duncan et al. (2017); Cheng et al. (2018), preconditioning AlRachid et al. (2018), and HFHR Li et al. (2022). We exclude Metropolis-Hastings and Hamiltonian Monte Carlo, as their acceptance-rejection steps require multiple score evaluations per step—which is too expensive compared to standard Langevin dynamics’ single evaluation.
After balancing interpretability, stability, and accuracy (Appendix F), we propose Fast Langevin Dynamics (FLD)—a variant of ULD defined by:
@@
where $\tau$ is the "time" of Langevin dynamics, $\mathbf{z}_{\tau}=(\mathbf{x}_{\tau},\mathbf{y}_{\tau})$ contains both the inpaint/known region, $\Gamma$ is the friction coefficient, $\mathbf{q}_{\tau}$ is the momentum. This dynamics is solved numerically using the FLD solver equation 78, which computes $\mathbf{z}_{\tau+\Delta\tau}$ analytically from $\mathbf{z}_{\tau}$. Details about FLD and its solver are discussed in Appendix F and Algorithm 4.
The FLD and its solver incorporates two key design features: (1) It introduces momentum $\mathbf{q}_{\tau}$ into the Langevin dynamics. Comparing with equation 1 reveals that equation 15 represents a time-averaged Langevin dynamics with decay rate $\Gamma$. This time averaging acts as momentum by incorporating memory of previous states, thereby accelerating convergence towards the stationary distribution. (2) To enhance the stability of FLD, we regularize its asymptotic behavior by decomposing the score function as $\mathbf{s}(\mathbf{z}_{\tau},t)=\mathbf{C}_{t}(\mathbf{z}_{\tau})-A_{t}\mathbf{z}_{\tau}$ in the FLD solver. The term $\mathbf{C}_{t}(\mathbf{z}_{\tau})$ is treated as constant over a time interval $[\tau,\tau+\Delta\tau]$, while $-A_{t}\mathbf{z}_{\tau}$ regulates the time step size, ensuring that $\mathbf{z}_{\tau+\Delta\tau}$ remains finite and stable, even for large $\Delta\tau$.
To understand how decomposition enhances stability, consider $A_{t}=(1-\bar{\alpha}_{t})^{-1}$. As $\Delta\tau\to\infty$, $\mathbf{z}_{\tau+\Delta\tau}$ remains finite and stable, following:
$\lim_{\Delta\tau\to\infty}\mathbf{z}_{\tau+\Delta\tau}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\hat{\mathbf{z}}_{0},1-\bar{\alpha}_{t}\right),$
where $\hat{\mathbf{z}}_{0}=\frac{\mathbf{z}_{\tau}+(1-\bar{\alpha}_{t})\mathbf{s}}{\sqrt{\bar{\alpha}_{t}}}$ is the Tweedie estimator for the clean image. This matches the forward diffusion process in equation 2, ensuring stability. Thus, large time steps can accelerate convergence without compromising output stability.
@@
Under the fast Langevin dynamics equation 15, the joint state $(\mathbf{z},\mathbf{q})$ has a stationary distribution given by
$(\mathbf{z},\mathbf{q})\;\sim\;p(\mathbf{z})\;\times\;\mathcal{N}(\mathbf{q}\mid\mathbf{0},\,\Gamma).$
Hence, $\mathbf{z}$ alone retains the same stationary distribution as the original Langevin dynamics equation 1. (Proof in Appendix H)
@@ 3.3 Rectified Flow Model Compatibility
We have introduced LanPaint using variance-preserving (VP) notation (Song et al., 2020b), corresponding to the forward diffusion process in equation 2. However, LanPaint is not limited to VP notation; it is general enough to apply to other mathematically equivalent diffusion frameworks, such as variance-exploding and rectified flow notations Liu et al. (2022b), by converting the score function into an eps-prediction or velocity prediction function, respectively. Detailed conversions among VP, variance-exploding, and rectified flow notations are provided in Appendix D.
@@ 4.1 Conditional Gaussian: Exactness of LanPaint
The task conditions on the y component to infer the x component. We compute the mean and covariance matrix of 50,000 samples and compare them with the ground truth distribution using KL divergence. The image shows the method comparisons across three plots: Ground Truth, KL Divergence vs. Diffusion Steps, and KL Divergence vs. Inner Iteration Steps. We adopt the same step size for Langevin-based methods (TFG and LanPaint).
Fig.2 demonstrates that LanPaint achieves near-zero KL divergence with the fewest diffusion steps and inner iteration steps, outperforming other methods. Fig.2 (right) also highlights that fast Langevin dynamics (FLD) alone, without BIG score, significantly accelerates convergence compared to the TFG method adopting the original Langevin dynamics.
A key observation is that heuristic linear inverse problem approaches (MCG, DPS, CoPaint, DDRM) cannot achieve KL divergence below 0.01 (dashed line) even with large number of steps or iterations, while exact conditional sampling methods (RePaint, TFG, LanPaint) succeed. This performance gap reflects their fundamental methodological difference: the former optimize heuristic objectives rather than the true distribution.
@@ 4.2 Mixture of Gaussian: Local Maxima Trapping
Figure 3: Local maxima trapping in inpainting $x$ given known $y$ using Euler sampler Karras et al. (2022). Red dots show unlikely samples trapped at local maxima of $p(x|y=1.55)$ (along the dashed line). Fewer diffusion steps (right $\rightarrow$ left) increase trapping—a key limitation of fast ODE sampler. Left panel shows LanPaint’s BiG score mitigates this issue. Methods perform 10 inner iterations/step (LanPaint, Langevin)
@@
The multi-modal Gaussian mixture distribution poses a significant challenge for inpainting with ODE samplers. As shown in Fig. 3, Langevin-based inpainting tends to concentrate samples at the "corners" of the distribution—local maxima of $p(x|y)$—despite their low joint likelihood $(x,y)$. This trapping phenomenon is not unique to Langevin methods; Fig. 4 shows that other approaches also produce samples clustered at the corners, where the distribution appears blurred.
Such trapping occurs due to insufficient information flow from the inpainted component $x$ to the observed $y$. During diffusion sampling, $x$ optimizes solely for $p(x|y)$, with no mechanism to penalize low $p(y|x)$. This motivates our BiG score equation 12, which propagates information from $x$ to $y$, steering samples away from low joint-likelihood regions, as shown in Fig.3.
Figure 4 compares sampling results and KL divergences across methods. Due to local maxima trapping, no method achieves zero KL divergence. However, LanPaint achieves significantly lower divergence than alternatives, demonstrating its effectiveness in mitigating trapping and producing accurate inpainting that closely matches the ground truth distribution.

### v3

@@ 4.1 Bidirectional Guided (BiG) Score
In RePaint and Langevin-based approaches, the inpainted region $\mathbf{x}_{t^{\prime}}$ at backward diffusion time $t^{\prime}$ is iterated towards high likelihood region of $p_{T-t^{\prime}}(\mathbf{x}\mid\mathbf{y})$, but the observed region $\mathbf{y}_{t^{\prime}}$ remains unaware of $\mathbf{x}_{t^{\prime}}$. This one-way dependency creates a critical flaw: if $\mathbf{x}_{t^{\prime}}$ enters a suboptimal region, $\mathbf{y}_{t^{\prime}}$ cannot receive corrective feedback, resulting in local maxima trapping of $\mathbf{x}_{t^{\prime}}$(Fig.3). To escape such local optima, we propose to jointly optimize $\mathbf{x}_{t^{\prime}}$ and $\mathbf{y}_{t^{\prime}}$ through bidirectional feedback: while $\mathbf{x}_{t^{\prime}}$ is refined by $\mathbf{y}_{t^{\prime}}$ (as in prior work), $\mathbf{y}_{t^{\prime}}$ is also updated under the guidance of $\mathbf{x}_{t^{\prime}}$ while preserving observed content.
To achieve this, we observe that Eq.7 is a special case of the following equivalent but more general form
$p_{t}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{o})\approx q_{\lambda,t}(\mathbf{x},\mathbf{y}\mid\mathbf{y}_{o})=\frac{1}{Z}\,p_{t}(\mathbf{x}\mid\mathbf{y})\,\tfrac{p_{t}(\mathbf{y}\mid\mathbf{y}_{o})^{1+\lambda}}{p_{t}(\mathbf{y})^{\lambda}},$ (11)
with $\mathbf{y}_{o}$ as the observed region of a given clean image, $\lambda>-1$ as the guidance scale and $Z$ is a normalizing constant. At $t=0$, the approximation ($\approx$) still becomes exact (==), which means that a sampled $\mathbf{x}$ from $q_{\lambda,t}$ is also a sample of the desired conditional distribution $p(\mathbf{x}\mid\mathbf{y}_{o})$ at $t=0$. The "=" holds at $t=0$ because $p_{0}(\mathbf{y}\mid\mathbf{y}_{o})=\delta(\mathbf{y}-\mathbf{y}_{o})$ is a delta distribution that remains invariant (up to normalization) when multiplied by other functions. Therefore $\tfrac{p_{t=0}(\mathbf{y}\mid\mathbf{y}_{o})^{1+\lambda}}{p_{t=0}(\mathbf{y})^{\lambda}}$ still represents $\delta(\mathbf{y}-\mathbf{y}_{o})$, enforcing that the sampled $\mathbf{y}_{t=0}$ equals $\mathbf{y}_{o}$.
Now the only problem that remains is how to let the diffusion model generate samples from $q_{\lambda,t}$ with Langevin dynamics. The $\mathbf{x}$ component of the score function of $q_{\lambda,t}$ is still $\mathbf{s}_{\mathbf{x}}$ in Eq.5, while the $\mathbf{y}$ component score can be approximated by the BiG score
@@
which is obtained by substituting $p_{t}(\mathbf{y})=\frac{p_{t}(\mathbf{x},\mathbf{y})}{p_{t}(\mathbf{x}|\mathbf{y})}$ into Eq.11 then discarding the unknown term $\nabla_{\mathbf{y}}\log p_{t}(\mathbf{x}|\mathbf{y})$. It successfully incorporates information of $\mathbf{x}_{t}$ as guidance for the $\mathbf{y}_{t}$ sampling process.
The BiG score’s behavior depends on $\lambda$: when $\lambda=-1$, it reduces to unconditional sampling; at $\lambda=0$, it matches RePaint-like inpainting; and for $\lambda>0$, it enhances inpainting by penalizing unconditional scores $\mathbf{s}_{\mathbf{y}}$. Larger $\lambda$ values strengthen the corrective feedback from $\mathbf{x}_{t}$, helping $\mathbf{y}_{t}$ to escape local optima more effectively.
@@
Though the unknown $\nabla_{\mathbf{y}}\log p_{t}(\mathbf{x}|\mathbf{y})$ term is discarded when deriving the BiG score, it still yields asymptotically exact conditional samples given partial observation, as this term is negligible (See Fig.9, Appendix.F) near $t=0$ compared to the score of $p_{t}(\mathbf{y}|\mathbf{y}_{o})$, as the following theorem states:
@@
Under the dynamics of Eq.13, the joint state $(\mathbf{x}_{t},\mathbf{y}_{t})$ converges to the distribution
$\mathbf{x}_{t},\,\mathbf{y}_{t}\;\sim\;\frac{1}{Z}\,p_{t}(\mathbf{x}\mid\mathbf{y})\,\tfrac{p_{t}(\mathbf{y}\mid\mathbf{y}_{o})^{1+\lambda}}{p_{t}(\mathbf{y})^{\lambda}}+\mathcal{O}(\sqrt{1-\bar{\alpha}_{t}}),$ (14)
where $Z$ is a normalizing constant. Consequently, at $t=0$, the marginal $\mathbf{x}_{0}\sim p_{0}(\mathbf{x}\mid\mathbf{y}_{o})$ is an exact conditional sample, provided the Langevin dynamics converges. (Proof in Appendix F)
@@ 4.2 Fast Langevin Dynamics (FLD)
Solving Langevin dynamics (Eq.13) is challenging: direct discretization requires trading step size against performance. Large steps accelerate convergence but introduce significant errors that yield noisy results, while small steps cause impractically slow convergence. We therefore seek an accelerated scheme with a stable solver, ensuring fast convergence to the stationary distribution while tolerating larger steps.
For fast convergence to the stationary distribution, existing approaches include Underdamped Langevin Dynamics (ULD) (Duncan et al., 2017; Cheng et al., 2018), preconditioning (AlRachid et al., 2018), and HFHR (Li et al., 2022). We exclude Metropolis-Hastings and Hamiltonian Monte Carlo, as their acceptance-rejection steps require multiple score evaluations per step—which is too expensive compared to standard Langevin dynamics’ single evaluation.
After balancing interpretability, stability, and accuracy (Appendix G), we propose Fast Langevin Dynamics (FLD)—a variant of ULD defined by:
@@
where $\tau$ is the "time" of Langevin dynamics, $\mathbf{z}_{\tau}=(\mathbf{x}_{\tau},\mathbf{y}_{\tau})$ contains both the inpainted/known region, $\Gamma$ is the friction coefficient, $\mathbf{q}_{\tau}$ is the momentum. This dynamics is solved numerically using the FLD solver Eq.115, which computes $\mathbf{z}_{\tau+\Delta\tau}$ analytically from $\mathbf{z}_{\tau}$. Details about FLD and its solver are discussed in Appendix G and Algorithm 4.
The FLD and its solver incorporates two key design features: (1) It introduces momentum $\mathbf{q}_{\tau}$ into the Langevin dynamics. Comparing with Eq.1 reveals that Eq.15 represents a time-averaged Langevin dynamics with decay rate $\Gamma$. This time averaging acts as momentum by incorporating memory of previous states, thereby accelerating convergence towards the stationary distribution. (2) We introduce a diffusion damping force when solving FLD numerically to enhance stability. The diffusion damping force is introduced by decomposing the score function as $\mathbf{s}(\mathbf{z}_{\tau},t)=\mathbf{C}_{t}(\mathbf{z}_{\tau})-A_{t}\mathbf{z}_{\tau}$ in the FLD solver. The term $\mathbf{C}_{t}(\mathbf{z}_{\tau})$ is treated as constant over a numerical time interval $\tau,\tau+\Delta\tau]$, while the diffusion damping force $A_{t}\mathbf{z}_{\tau}$ serves as a regularization inspired by the forward diffusion process, ensuring that $\mathbf{z}_{\tau+\Delta\tau}$ remains finite and stable, even for large $\Delta\tau$.
To understand how the diffusion damping force is related to the diffusion model and enhances stability, consider $A_{t}=(1-\bar{\alpha}_{t})^{-1}$. As $\Delta\tau\to\infty$, $\mathbf{z}_{\tau+\Delta\tau}$ remains finite and stable, following:
$\lim_{\Delta\tau\to\infty}\mathbf{z}_{\tau+\Delta\tau}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\hat{\mathbf{z}}_{0},1-\bar{\alpha}_{t}\right),$ (16)
where $\hat{\mathbf{z}}_{0}=\frac{\mathbf{z}_{\tau}+(1-\bar{\alpha}_{t})\mathbf{s}}{\sqrt{\bar{\alpha}_{t}}}$ is the Tweedie estimator for the clean image. This matches the forward diffusion process in Eq.2, ensuring stability. Thus, large time steps can accelerate convergence without compromising output stability.
@@
Under the fast Langevin dynamics Eq.15, the joint state $(\mathbf{z},\mathbf{q})$ has a stationary distribution given by
$(\mathbf{z},\mathbf{q})\;\sim\;p(\mathbf{z})\;\times\;\mathcal{N}(\mathbf{q}\mid\mathbf{0},\,\Gamma).$ (17)
Hence, $\mathbf{z}$ alone retains the same stationary distribution as the original Langevin dynamics Eq.1. (Proof in Appendix H)
@@ 4.3 Rectified Flow Model Compatibility
We have introduced LanPaint using variance-preserving (VP) notation (Song et al., 2020c), corresponding to the forward diffusion process in Eq.2. However, LanPaint is not limited to VP notation; it is general enough to apply to other mathematically equivalent diffusion frameworks, such as variance-exploding and rectified flow notations (Liu et al., 2022b), by converting the score function into an eps-prediction or velocity prediction function, respectively. Detailed conversions among VP, variance-exploding, and rectified flow notations are provided in Appendix E.
@@
4.4 Generalization to Arbitrary Dimensions
LanPaint’s core formulations, the BiG score (Eq. 12) and FLD (Eq. 15)—are dimension-agnostic, operating on joint scores $\mathbf{s}(\mathbf{x},\mathbf{y},t)$ for arbitrary-dimensional tensors $\mathbf{z}=(\mathbf{x},\mathbf{y})\in\mathbb{R}^{d}$ ($d\geq 1$) without 2D-specific assumptions. This enables exact conditional sampling $\mathbf{x}\sim p(\mathbf{x}\mid\mathbf{y})$ across modalities (e.g., 1D audio sequences, 2D images, 3D volumes, spatio-temporal data such as video). As an example, we extend LanPaint to video inpainting, treating video as a tensor $\mathbf{z}\in\mathbb{R}^{F\times H\times W}$ (stacked frames ${\mathbf{z}^{(f)}}\in\mathbb{R}^{H\times W},f=1,...,F$) with static masks in Sec.5.6.
@@ 5.1 Conditional Gaussian: Exactness of LanPaint
The task conditions on the y component to infer the x component. We compute the mean and covariance matrix of 50,000 samples and compare them with the ground truth distribution using KL divergence. Fig.2 shows the method comparisons across three plots: Ground Truth, KL Divergence vs. Diffusion Steps, and KL Divergence vs. Inner Iteration Steps. We adopt the same step size for Langevin-based methods (TFG and LanPaint).
Fig.2 also demonstrates that LanPaint achieves near-zero KL divergence with the fewest diffusion steps and inner iteration steps, outperforming other methods. Fig.2 (right) also highlights that fast Langevin dynamics (FLD) alone, without BiG score, significantly accelerates convergence compared to the TFG method adopting the original Langevin dynamics.
A key observation is that heuristic linear inverse problem approaches (MCG, DPS, CoPaint, DDRM) cannot achieve KL divergence below 0.01 (dashed line) even with a large number of steps or iterations, while exact conditional sampling methods (RePaint, TFG, LanPaint) succeed. This performance gap reflects their fundamental methodological difference: the former optimize heuristic objectives rather than the true distribution.
@@ 5.2 Mixture of Gaussian: Local Maxima Trapping
Figure 3: Local maxima trapping in inpainting $x$ given known $y$ using Euler sampler (Karras et al., 2022). Red dots show unlikely samples trapped at local maxima of $p(x|y=1.55)$ (along the dashed line). Fewer diffusion steps (right $\rightarrow$ left) increase trapping—a key limitation of fast ODE sampler. Left panel shows LanPaint’s BiG score mitigates this issue. Methods perform 10 inner iterations/step (LanPaint, Langevin)
The multi-modal Gaussian mixture distribution poses a significant challenge for inpainting with ODE samplers. As shown in Fig.3, Langevin-based inpainting tends to concentrate samples at the "corners" of the distribution—local maxima of p(x|y)p(x|y)—despite their low joint likelihood $(x,y)$. This trapping phenomenon is not unique to Langevin methods; Fig.4 shows that other approaches also produce samples clustered at the corners, where the distribution appears blurred.
@@
Such trapping occurs due to insufficient information flow from the inpainted component $x$ to the observed $y$. During diffusion sampling, $x$ optimizes solely for $p(x|y)$, with no mechanism to penalize low $p(y|x)$. This motivates our BiG score Eq.12, which propagates information from $x$ to $y$, steering samples away from low joint-likelihood regions, as shown in Fig.3.
Fig.4 compares sampling results and KL divergences across methods. Due to local maxima trapping, no method achieves zero KL divergence. However, LanPaint achieves significantly lower divergence than alternatives, demonstrating its effectiveness in mitigating trapping and producing accurate inpainting that closely matches the ground truth distribution.